Find the eigenvalues and eigenfunctions of the following Sturm-Liouville probelm (e^(2x)y')' + e^(2x) (λ + 1)y = 0; y(0) = 0 = y(π). 


1
Expert's answer
2021-10-05T17:06:06-0400

"(e^{2x}y')'+e^{2x}(\lambda +1)y=0\\
e^{2x}y''+2e^{2x}y''+e^{2x}(\lambda +1)y=0\\
\text{Divide through by } e^{2x}\\
y''+2y'+(\lambda + 1)=0\\
\text{We can only have a solution when } \lambda >0\\
m^2+2m+(\lambda + 1)=0\\
m=-1\pm\sqrt{-\lambda}\\
y(x)=e^{-1}[C_1\cos(x\sqrt{\lambda})+C_2\sin(x\sqrt{\lambda})]\\
\text{Apply the boundary conditions}\\
y(0)=e^{-1}C_1 \implies C_1=0\\
y(\pi)=e^{-1}[C_2\sin(\pi \sqrt{\lambda}))\\
C_2 \neq 0\\
\implies \sin(\pi \sqrt{\lambda})=0\\
\implies \pi \sqrt{\lambda}=n\pi ~~~n=1,2,3, \cdots\\
\implies \lambda_n =n^2
\text{This is the Eigen values}\\
y_n(x)=e^{-1}[C_n\sin(nx)]\\
\text{This is the Eigen Functions}"